For the consecutive reactions A ‹ B‹C, k1=0.35/h, k2=0.13/h, CA0=4 lbmol/ft concentration of B is maximum.and CB0=0, CC0=0. Find the time when the concentration of B is maximum.

  1. 7.2 hrs                  b.             6.5 hrs                   c.            4.6 hrs                  d.            3.9 hrs
  1. Based from the preceding problem, what is the maximum concentration of B in lbmols/ft3 if the reactor used is a single CSTR?

a.              1.85                     b.               2.01                      c.               1.22                    d.              2.32
Answer #1

Solutions k9. k A B C ki = 0.35/h kg = 0.13/h CAO = 4 lbmor/ft (Boa Cco=0 time when the Concentration of bis maximum as is gi

Thank you.... Please upvote